Light travels at a constant speed of 3.00x10^8 m/s, whereas sound travels through the air at a constant speed of 343 m/s. How long does it take for light to travel from a lightning strike to an observer 4.10 km away?
1
Expert's answer
2020-10-27T11:15:09-0400

We can find the time that the light takes to travel from a lightning strike to an observer from the formula:


"t=\dfrac{s}{v}=\dfrac{4.10\cdot10^3\ m}{3.0\cdot 10^8\ \dfrac{m}{s}}=1.36\cdot 10^{-5}\ s."

Answer:

"t=1.36\cdot 10^{-5}\ s."